# Correlation between the first-reaction time and the acquired boundary local time

### v1

@@ I Introduction
The stochastic motion of particles and resulting encounters with their partners lie at the heart of numerous physical, chemical, and biological processes [1, 2, 3, 4, 5]. For diffusive particles inside a bounded domain $\Omega$ with Dirichlet boundary condition, the statistical characterization of these encounters through the distribution of first-passage times (FPTs) has been a central theme in statistical physics for decades [6, 7, 2, 8, 5, 9, 10, 3, 11, 13, 12, 14, 15, 4], where single or multiple targets can be either inside $\Omega$, or distributed on the boundary $\partial\Omega$}. In this classical framework, a target is considered perfectly absorbing, meaning the trajectory of a particle terminates upon the first encounter with the target. The FPT to such a target has been extensively analyzed, yielding profound insights into diffusion-limited kinetics [16, 17]. However, the first arrival of the particle on the reactive boundary does not necessarily achieve a successful absorption or reaction event. In most realistic scenarios, targets are not perfect sinks [4, 18, 19]. Examples range from trapping of reactants on catalytic surfaces [20, 21, 22] and binding of ligands to cellular receptors [23, 24] to the search of proteins for specific DNA sites [25, 26]. Microscopically, a molecule may interact with a reactive site multiple times due to insufficient overlapping between atom orbitals to form chemical bonds, thus being reflected back into the bulk after each non-reactive collision [27, 28]. Such reflected Brownian dynamics result in a random number of encounters with reactive sites before the successful absorption or reaction event occurs at the end. Therefore, the FPT can be expanded beyond the idealized scenario of a perfectly absorbing boundary to a more general concept of the first-reaction time (FRT) on a partially reactive boundary. In this setting, the propensity of the surface to react is characterized by a constant reactivity $\kappa$, while the ratio $q=\kappa/D$ quantifies the balance between diffusive and reactive fluxes on the boundary [21, 29]. The limit $q\to\infty$ recovers the perfect absorption such that the FRT is reduced to the FPT, while $q=0$ describes a perfectly reflecting, inert surface and the FRT turns out to be infinite.
@@
where $D$ is the diffusion coefficient, $|\bm{x}-\Gamma|$ is the Euclidean distance between a point $\bm{x}$ and the reactive region $\Gamma$, and $\Theta(z)$ is the Heaviside step function: $\Theta(z)=1$ for $z>0$ and 0 otherwise. The BLT is used to construct the encounter-based approach to diffusion-controlled reactions [33], allowing one to consider different surface reaction mechanisms by imposing suitable stopping conditions [34, 35, 36, 37, 38, 39]. For instance, the reaction event on a partially reactive boundary occurs when $\ell_{t}$ exceeds an independent, exponentially distributed random threshold $\hat{\ell}$, satisfying $\mathbb{P}\{\hat{\ell}>\ell\}=e^{-q\ell}$. As a consequence, the first-reaction time on $\Gamma$ can be defined as:
@@
This probabilistic formulation explicitly connects the macroscopic parameter $q$ to a microscopic stopping condition for the BLT. Still, the acquired BLT does not uniquely constrain the FRT. Even though two diffusive particles starting from the same location may acquire the same BLT, their FRTs can be substantially different. In other words, one cannot infer the FRT from the BLT, nor vice versa; in fact, these random variables provide complementary information on the reaction event. This raises a fundamental and yet unexplored question: How statistically correlated are the first-reaction time $\tau$ and the boundary local time $\ell_{\tau}$ acquired until the instance of reaction on $\Gamma$? In particular, the correlation coefficient $C_{q}=\mathrm{Corr}(\tau,\ell_{\tau})$ that measures the degree to which reaction time predicts boundary exploration (and vice versa), is crucial for deeper understanding of the diffusion-reaction pathways.
@@ II Theoretical results
where $\partial_{n}$ is the normal derivative oriented outward from the domain $\Omega$. The spectrum of this problem is known to be discrete [40], and the eigenpairs $\{\mu_{k}^{(p)},V_{k}^{(p)}\}$ are enumerated by $k=0,1,2,\ldots$ to order the eigenvalues $\mu_{k}^{(p)}$ into an increasing sequence. The moment of the FCT can be found from the moment-generating function $\tilde{U}(\ell,p|\bm{x}_{0})$:
@@
According to Eq. (3), the first-reaction time is the FCT of the random threshold $\hat{\ell}$, i.e., $\tau=\mathcal{T}_{\hat{\ell}}$. As a consequence, its probability density $H_{q}(t|\bm{x}_{0})$ can be obtained by averaging $U(\ell,t|\bm{x}_{0})$ over the exponential distribution of the threshold $\hat{\ell}$:
$H_{q}(t|\bm{x}_{0})=\langle U(\hat{\ell},t|\bm{x}_{0})\rangle=\int\limits_{0}^{\infty}\mathrm{d}\ell\ qe^{-q\ell}\ U(\ell,t|\bm{x}_{0}).$ (8)
Moreover, this formula indicates that
$\mathcal{P}_{q}(\ell,t|\bm{x}_{0})=qe^{-q\ell}\ U(\ell,t|\bm{x}_{0})\,$, (9)
is actually the joint probability density of $\ell_{\tau}$ and $\tau$. This conceptually important extension of $U(\ell,t|\bm{x}_{0})$ was unnoticed in previous works, and presents one of the main results of this paper. In particular, $H_{q}(t|\bm{x}_{0})$ is the marginal probability density of $\tau$, whereas $qe^{-q\ell}$ is the marginal probability density of $\ell_{\tau}=\hat{\ell}$.
For any integer $m>0$, one can express the $m$th moment of the first-reaction time as:
$\langle\tau^{m}\rangle$ $=\int\limits_{0}^{\infty}\mathrm{d}t\ t^{m}H_{q}(t|\bm{x}_{0})=(-1)^{m}\left(\partial_{p}^{m}\tilde{H}_{q}(p|\bm{x}_{0})\right)_{p=0}$
$=(-1)^{m}\int\limits_{0}^{\infty}\mathrm{d}\ell\ qe^{-q\ell}\left(\partial^{m}_{p}\tilde{U}(\ell,p|\bm{x}_{0})\right)_{p=0}$
$=\int\limits_{0}^{\infty}\mathrm{d}\ell\ qe^{-q\ell}\langle\mathcal{T}^{m}_{\ell}\rangle,$ (10)
where $\tilde{H}_{q}(p|\bm{x}_{0})$ is the Laplace transform of $H_{q}(t|\bm{x}_{0})$. An alternative calculation can be realized by the following recurrence relations for $m\geqslant 1$:
@@
with $\langle\tau^{0}\rangle=1$.
In turn, the $m$th moment of the acquired boundary local time $\ell_{\tau}$ is universal and independent of the domain $\Omega$:
$\langle\ell_{\tau}^{m}\rangle=\int\limits_{0}^{\infty}\mathrm{d}\ell\ \ell^{m}\int\limits_{0}^{\infty}\mathrm{d}t\ qe^{-q\ell}\ U(\ell,t|\bm{x}_{0})=\frac{m!}{q^{m}},$ (12)
where we used the normalization of the probability density $U(\ell,t|\bm{x}_{0})$, such that $\int_{0}^{\infty}\mathrm{d}t\ U(\ell,t|\bm{x}_{0})=1$. In particular, one gets the variance of $\ell_{\tau}$ as:
@@
One also computes the crossing average as:
@@ II.1 Asymptotic behavior
As the starting point does not matter in the leading order, i.e., $\langle\tau\rangle\approx$ const, one has
@@ II.2 Special case for the starting point
and the correlation coefficient in this case becomes:
@@ II.3 Example of a ball
To illustrate the above results, we consider diffusion in a ball $\Omega=\{\bm{x}\in\mathbb{R}^{3}:|\bm{x}|<R\}$ of radius $R$, with a partially reactive boundary ${\partial\Omega}=\Gamma$ (i.e., ${\partial\Omega}_{N}=\emptyset$), for which the Steklov eigenvalues and eigenfunctions are known explicitly (see, e.g., [33, 41]). The rotational symmetry implies that only the rotationally invariant eigenfunction $V_{0}^{(p)}(\bm{x})=\frac{1}{\sqrt{|\Gamma|}}\frac{i_{0}(\alpha|\bm{x}|)}{i_{0}(\alpha R)}$ contributes to Eq. (5), yielding
@@
where $\alpha=\sqrt{p/D}$, $\mu_{0}^{(p)}=\alpha i_{0}^{\prime}(\alpha R)/i_{0}(\alpha R)$, $i_{0}(z)=\sinh(z)/z$, and $i_{0}^{\prime}(z)=\mathrm{d}i_{0}(z)/\mathrm{d}z$. For a starting point $|\bm{x}_{0}|=r$ inside the ball, one retrieves the Laplace transform of the FRT probability density:
@@
One can easily check the asymptotic behavior described in Sec. II.1: $C_{q}(r)\to 1$ as $q\to 0$ and $C_{q}(r)\to 0$ as $q\to\infty$. Moreover, we quantify now how fast or slow these limits are achieved; in particular, as $q\to\infty$, one gets $C_{q}(r)\propto q^{-1}$ for any $r<R$ and $C_{q}(R)\propto q^{-1/2}$, as we saw in Sec. II.4.
In Appendix A.1, we also derive the correlation coefficient for the starting point uniformly distributed inside the ball (we use the symbol $\circ$ to refer to this setting):
@@ II.4 Concentric spherical shell $(d\geq 2)$
These asymptotic results clearly show that the FRT and the BLT are strongly correlated in the small-target limit with moderate reactivity $q$ (such that $qR\ll 1$). We stress that the reactivity itself does not need to be small: one needs $q\ll q_{\mathrm{max}}$, with the maximal value $q_{\mathrm{max}}=1/R$ being large as $R\to 0$. While Eqs. (41, 42) were derived for a particular geometry of a spherical shell, we expect that the above conclusion remains valid in general.
@@ III Numerical results
In this section, we employ Monte Carlo simulations to provide complementary insights into correlations between the FRT and the BLT. Sections III.1 and III.2 present the simulation algorithm and its validation, whereas Sections III.3, III.4, III.5 summarize our main numerical results for restricted diffusion in disordered media.
To investigate the role of geometric disorder, we consider domains containing interior inert obstacles (Fig. 1). We study three packing configurations: (a) a regular arrangement of identical disks of radius $R$ inside the unit disk, with the central disk being reactive and all the others being reflecting; (b) a random packing of identical non-overlapping disks of radius $R$ inside the unit disk; (c) a regular arrangement of identical balls of radius $R$ inside the unit ball, with the central ball being reactive.
@@ III.1 Description of Monte Carlo simulations
Figure 1: Computational domains with interior obstacles of the same radius $R$: the central one (in blue) is the reactive target, while the others are inert. (a) Regular square-lattice packing of $m\times m$ identical disks inside the unit disk, with $m=5$. The shown configuration has the maximal radius $R_{\mathrm{max}}$ such that all disks touch each other compactly without overlapping. (b) Random packing of 25 disks with the maximal radius $R_{\mathrm{max}}$ that corresponds to a fixed area occupation of 50%. The coordinates of all disks are provided in Appendix B. (c) Regular cubic-lattice packing of $m\times m\times m$ identical balls inside the unit ball with $m=3$, and the maximal radius $R_{\mathrm{max}}$.
@@ III.2 Validation
Figure 2: Correlation coefficient $C_{q}(\circ)$ for (a) a unit ball and (b) a unit disk ($R=1$). Comparison between Monte Carlo results (MC, blue dots) and analytical expressions (39, 40) (orange curves). Numerical statistics are obtained from $N=10^{6}$ particles, with $D=1$. Green dashed lines indicate the asymptotic behaviors $C_{q}(\circ)\approx 5\sqrt{7/13}/(qR)$ for $d=3$ and $C_{q}(\circ)\approx 4\sqrt{3/5}/(qR)$ for $d=2$ in the limit $q\to\infty$.
@@
Our Monte Carlo method is validated by comparing the correlation coefficients $C_{q}(\circ)$ obtained from simulations with the exact analytical expressions (39, 40) for uniform starting points inside the unit ball ($d=3$) or the unit disk ($d=2$). As shown in Fig. 2(a), the numerical results (blue dots) exhibit excellent agreement with Eq. (39) (orange curve) across a broad range of $q$, from $10^{-2}$ to $10^{3}$. In the limit $q\to 0$, one gets $C_{q}\to 1$, indicating that the FRT $\tau$ and the BLT $\ell_{\tau}$ become fully correlated. As $q$ increases, $C_{q}(\circ)$ decreases monotonically. In the large reactivity limit $q\to\infty$, a leading-order expansion of Eq. (39) yields the asymptotic behavior $C_{q}(\circ)\propto 1/q$. The same behavior is observed for the unit disk as shown in Fig. 2(b), where the simulation data match the exact solution given by Eq. (40).
@@
Figure 3: Joint histograms of the FRT $\tau$ and the BLT $\ell_{\tau}$ for $q=10^{-2},10^{-1},10^{0},10^{1}$ obtained by Monte Carlo simulations with $N=10^{6}$ particles, where the starting points are uniform inside the unit ball. Both $\tau$ and $\ell_{\tau}$ are normalized by their respective mean values. In each panel, the scatter of blue dots represents individual Monte Carlo results of the pair $(\tau,\ell_{\tau})$, whose density approximated the underlying joint probability density $\mathcal{P}_{q}(\ell,t|\circ)$ from Eq. (9). The marginal distributions of $\tau/\langle\tau\rangle$ and $\ell_{\tau}/\langle\ell_{\tau}\rangle=q\ell_{\tau}$ are shown above and right of each joint plot.
@@
Beyond the correlation coefficient itself, the joint histogram of $\tau$ and $\ell_{\tau}$ offers a more direct visual representation of their statistical coupling. Figure 3 presents such joint histograms for uniform starting points inside the unit ball at different reactivities. Both axes are normalized by their mean values given by Eq. (48a) and Eq. (12), respectively. The marginal distributions of $\tau$ and $\ell_{\tau}$ exhibit similar exponential decay across all $q$. For large $q$, the pairs $(\tau,\ell_{\tau})$ are widely scattered, reflecting weak correlations. In turn, for smaller reactivity, pairs concentrate along a straight line $\ell_{\tau}/\langle\ell_{\tau}\rangle\approx\tau/\langle\tau\rangle$, indicating the strong linear relation between $\tau$ and $\ell_{\tau}$ anticipated from the limit $C_{q}\to 1$ as $q\to 0$.

### v2

@@ I Introduction
The stochastic motion of particles and resulting encounters with their partners lie at the heart of numerous physical, chemical, and biological processes [1, 2, 3, 4, 5]. For diffusive particles inside a bounded domain $\Omega$ with Dirichlet boundary condition, the statistical characterization of these encounters through the distribution of first-passage times (FPTs) has been a central theme in statistical physics for decades [6, 7, 2, 8, 5, 9, 10, 3, 11, 13, 12, 14, 15, 4], where single or multiple targets can be either inside $\Omega$, or distributed on the boundary $\partial\Omega$}. In this classical framework, a target is considered perfectly absorbing: the trajectory of a particle terminates upon the first encounter with the target. The FPT to such a target has been extensively analyzed, yielding profound insights into diffusion-limited kinetics [16, 17]. However, the first arrival of the particle on the reactive boundary does not necessarily achieve a successful absorption or reaction event. In most realistic scenarios, targets are not perfect sinks [4, 18, 19]. Examples range from trapping of reactants on catalytic surfaces [20, 21, 22] and binding of ligands to cellular receptors [23, 24] to the search of proteins for specific DNA sites [25, 26]. Microscopically, a molecule may interact with a reactive site multiple times due to insufficient overlapping between atom orbitals to form chemical bonds, thus being reflected back into the bulk after each non-reactive collision [27, 28]. Such reflected Brownian dynamics result in a random number of encounters with reactive sites before the successful absorption or reaction event occurs at the end. Therefore, the FPT can be expanded beyond the idealized scenario of a perfectly absorbing boundary to a more general concept of the first-reaction time (FRT) on a partially reactive boundary. In the most common setting, the propensity of the surface to react is characterized by a constant reactivity $\kappa$, while the ratio $q=\kappa/D$ quantifies the balance between diffusive and reactive fluxes on the boundary [21, 29]. The limit $q\to\infty$ recovers the perfect absorption such that the FRT is reduced to the FPT, while $q=0$ describes a perfectly reflecting, inert surface and the FRT turns out to be infinite.
@@
where $D$ is the diffusion coefficient, $|\bm{x}-\Gamma|$ is the Euclidean distance between a point $\bm{x}$ and the reactive region $\Gamma$, and $\Theta(z)$ is the Heaviside step function: $\Theta(z)=1$ for $z>0$ and 0 otherwise.
The BLT is used to construct the encounter-based approach to diffusion-controlled reactions [33], allowing one to describe different surface reaction mechanisms by imposing suitable stopping conditions [34, 35, 36, 37, 38, 39]. In this paper, we consider a broad class of surface reactions, in which the reaction event occurs after a sufficient number of encounters of the particle with the reactive region $\Gamma$. More formally, we introduce an independent random variable $\hat{\ell}$ with some probability density $\psi(\ell)$ and postulate that the first-reaction time on $\Gamma$ is the first instance when the BLT $\ell_{t}$ exceeds the threshold $\hat{\ell}$:
@@
In this way, the BLT $\ell_{t}$ incorporates the diffusive dynamics in $\Omega$, whereas the threshold $\hat{\ell}$ describes the chemical kinetics on the reactive region. For instance, the most common scenario of a partially reactive boundary with a constant reactivity $\kappa=qD$ corresponds to the exponential threshold with the density $\psi(\ell)=qe^{-q\ell}$ (see [33] for details).
This probabilistic formulation explicitly connects the macroscopic parameter $q$ to a microscopic stopping condition for the BLT. Still, the acquired BLT does not uniquely constrain the FRT. Even though two diffusive particles starting from the same location may acquire the same BLT, their FRTs can be substantially different. In other words, one cannot infer the FRT from the BLT, nor vice versa; in fact, these random variables provide complementary information on the reaction event. This raises a fundamental and yet unexplored question: How statistically correlated are the first-reaction time $\tau$ and the boundary local time $\ell_{\tau}$ accumulated until the instance of reaction on $\Gamma$? In particular, the correlation coefficient $C_{q}=\mathrm{Corr}(\tau,\ell_{\tau})$ that measures the degree to which reaction time predicts boundary exploration (and vice versa), is crucial for deeper understanding of the diffusion-reaction pathways.
@@ II Theoretical results
where $\partial_{n}$ is the normal derivative oriented outward from the domain $\Omega$. The spectrum of this problem is known to be discrete [40], and the eigenpairs $\{\mu_{k}^{(p)},V_{k}^{(p)}\}$ are enumerated by $k=0,1,2,\ldots$ to order the eigenvalues $\mu_{k}^{(p)}$ into an increasing sequence. The $m$th-order moment of the FCT can be found from the moment-generating function $\tilde{U}(\ell,p|\bm{x}_{0})$:
@@
Let us now show that the probability density $U(\ell,t|\bm{x}_{0})$ plays the key role for understanding correlations between the first-reaction time $\tau$ and the BLT $\ell_{\tau}$ prior to the reaction event. According to Eq. (3), the first-reaction time is the FCT of the random threshold $\hat{\ell}$, i.e., $\tau=\mathcal{T}_{\hat{\ell}}$ and thus $\ell_{\tau}=\hat{\ell}$. As a consequence, the joint probability density of $\ell_{\tau}$ and $\tau$ is
$\mathcal{P}(\ell,t|\bm{x}_{0})=\psi(\ell)\,U(\ell,t|\bm{x}_{0}).$ (8)
In fact, once the random threshold $\hat{\ell}$ is selected with the density $\psi(\ell)$, the first-reaction time is chosen with the density $U(\ell,t|\bm{x}_{0})$. The product in Eq. (8) simply reflects that the threshold was supposed to be independent of the diffusive dynamics (and thus of the confining domain); in this way, the threshold characterizes exclusively the chemical kinetics on the reactive region. This conceptually important extension of $U(\ell,t|\bm{x}_{0})$ was unnoticed in previous works, and presents one of the main results of this paper. Note that the (marginal) probability density of $\tau$ is obtained by integrating $\mathcal{P}(\ell,t|\bm{x}_{0})$ over $\ell$, whereas the (marginal) probability density of $\ell_{\tau}$ is obtained by integrating over $t$, which expectedly yields $\psi(\ell)$ due to the normalization of $U(\ell,t|\bm{x}_{0})$:
$\int\limits_{0}^{\infty}\mathrm{d}t\ U(\ell,t|\bm{x}_{0})=1.$ (9)
In the following, we focus on the most common scenario of a partially reactive region with a constant reactivity $\kappa=qD$, for which $\psi(\ell)=qe^{-q\ell}$. In this case, the probability density of the first-reaction time, denoted as $H_{q}(t|\bm{x}_{0})$, turns out to be proportional to the Laplace transform of $U(\ell,t|\bm{x}_{0})$ with respect to $\ell$:
$H_{q}(t|\bm{x}_{0})=\int\limits_{0}^{\infty}\mathrm{d}\ell\ {\mathcal{P}}(\ell,t|\bm{x}_{0})=\int\limits_{0}^{\infty}\mathrm{d}\ell\ qe^{-q\ell}\ U(\ell,t|\bm{x}_{0}).$ (10)
For any integer $m>0$, one can express the $m$th-order moment of the first-reaction time as:
$\langle\tau^{m}\rangle$ $=\int\limits_{0}^{\infty}\mathrm{d}\ell\int\limits_{0}^{\infty}\mathrm{d}t\ t^{m}\ {\mathcal{P}}(\ell,t|\bm{x}_{0})=\int\limits_{0}^{\infty}\mathrm{d}\ell\ qe^{-q\ell}\langle\mathcal{T}^{m}_{\ell}\rangle.$ (11)
An alternative calculation can be realized by the following recurrence relations for $m\geqslant 1$:
@@
with $\langle\tau^{0}\rangle=1$. As discussed in [33], the Robin boundary condition (12a) is tightly related to the chosen exponential form of $\psi(\ell)$; in particular, another probability law for the threshold $\hat{\ell}$ would result in a different boundary condition.
In turn, the $m$th-order moment of the acquired boundary local time $\ell_{\tau}$ is universal and independent of the domain $\Omega$:
$\langle\ell_{\tau}^{m}\rangle=\int\limits_{0}^{\infty}\mathrm{d}\ell\ \ell^{m}\int\limits_{0}^{\infty}\mathrm{d}t\ {\mathcal{P}}(\ell,t|\bm{x}_{0})=\frac{m!}{q^{m}},$ (13)
where we used again the normalization (9) of the probability density $U(\ell,t|\bm{x}_{0})$. This is simply the $m$th-order moment of the chosen exponentially distributed threshold $\hat{\ell}$. In particular, one gets the variance of $\ell_{\tau}$ as:
@@
We also compute the crossing average as:
@@ II.1 Asymptotic behavior
As the starting point does not matter in the leading order in this regime, one has $\langle\tau\rangle\approx$ const, and thus
@@ II.2 Special case for the starting point
and the correlation coefficient in this case becomes
@@ II.3 Example of a ball
To illustrate the above results, we consider diffusion in a ball $\Omega=\{\bm{x}\in\mathbb{R}^{3}:|\bm{x}|<R\}$ of radius $R$, with a partially reactive boundary ${\partial\Omega}=\Gamma$ (here ${\partial\Omega}_{N}=\emptyset$), for which the Steklov eigenvalues and eigenfunctions are known explicitly (see, e.g., [33, 41]). The rotational symmetry implies that only the rotationally invariant eigenfunction $V_{0}^{(p)}(\bm{x})=\frac{1}{\sqrt{|\Gamma|}}\frac{i_{0}(\alpha|\bm{x}|)}{i_{0}(\alpha R)}$ contributes to Eq. (5), yielding
@@
where $\alpha=\sqrt{p/D}$, $\mu_{0}^{(p)}=\alpha i_{0}^{\prime}(\alpha R)/i_{0}(\alpha R)$, $i_{0}(z)=\sinh(z)/z$ is the modified spherical Bessel function of the first kind and zero order, and $i_{0}^{\prime}(z)=\mathrm{d}i_{0}(z)/\mathrm{d}z$. For a starting point $|\bm{x}_{0}|=r$ inside the ball, one retrieves the Laplace transform of the FRT probability density:
@@
One can easily check the asymptotic behavior described in Sec. II.1: $C_{q}(r)\to 1$ as $q\to 0$ and $C_{q}(r)\to 0$ as $q\to\infty$. Moreover, we quantify now how fast or slow these limits are achieved; in particular, as $q\to\infty$, one gets $C_{q}(r)\propto q^{-1}$ for any $r<R$ and $C_{q}(R)\propto q^{-1/2}$, as we saw in Sec. II.2. A similar computation for the disk of radius $R$ yields (see Appendix A.2):
$C_{q}(r)=\frac{1}{\sqrt{1+qR/2+(qR)^{2}\left(1-(r/R)^{4}\right)/8}}.$ (40)
In Appendix A.1, we also derive the correlation coefficient for the case when the starting point is uniformly distributed inside the ball (we use the symbol $\circ$ to refer to this setting):
@@ II.4 Concentric spherical shell $(d\geq 2)$
These asymptotic results clearly show that the FRT and the BLT are strongly correlated in the small-target limit with moderate reactivity $q$ (such that $qR\ll 1$). We stress that the reactivity itself does not need to be small: one needs $q\ll q_{\mathrm{max}}$, with the maximal value $q_{\mathrm{max}}=1/R$ being large as $R\to 0$. While Eqs. (43, 44) were derived for a particular geometry of a spherical shell, we expect that the above conclusion remains valid in general.
@@ III Numerical results
In the previous section, we have shown that the probability density $U(\ell,t|\bm{x}_{0})$ of the first-crossing time $\mathcal{T}_{\ell}$ determines the joint probability density $\mathcal{P}(\ell,t|\bm{x}_{0})$ and controls correlations between the first-reaction time $\tau$ and the acquired boundary local time $\ell_{\tau}$. In particular, the theoretical framework based on the generalized Steklov problem gives access to the moment-generating function $\tilde{U}(\ell,p|\bm{x}_{0})$ of $\mathcal{T}_{\ell}$ via its spectral expansion. We illustrated its efficient application for rotation-invariant domains. In general, however, the Steklov eigenfunctions are not known explicitly, and one has to resort to numerical techniques such as a finite-element method [42, 43] or a boundary-integral method [44]. In the particular case when the reactive region $\Gamma$ is small, one can employ matched asymptotic techniques to get explicit but approximate results [35, 36, 45, 46].
In this section, we choose a different strategy and employ Monte Carlo simulations to provide complementary insights into correlations between the FRT and the BLT. Sections III.1 and III.2 present the simulation algorithm and its validation, whereas Sections III.3, III.4, III.5 summarize our main numerical results for restricted diffusion in disordered media.
To investigate the role of geometric disorder, we consider domains containing interior inert obstacles (Fig. 1). We study three packing configurations: (a) a regular arrangement of identical disks of radius $R$ inside the unit disk; (b) a random packing of identical non-overlapping disks of radius $R$ inside the unit disk; (c) a regular arrangement of identical balls of radius $R$ inside the unit ball, with the central ball being reactive. In all three cases, the central disk (ball) is partially reactive, whereas all the others are reflecting (inert obstacles).
@@ III.1 Description of Monte Carlo simulations
Figure 1: Computational domains with interior obstacles of the same radius $R$: the central one (in blue) is the reactive target, while the others are inert. (a) Regular square-lattice packing of $m\times m$ identical disks inside the unit disk, with $m=5$. The shown configuration has $R=0.618\,R_{\mathrm{max}}$, with the maximal radius $R_{\mathrm{max}}=L/(\sqrt{2}(m-1)+1)\approx 0.15$. (b) Random packing of 25 disks with the same $R$. The coordinates of all disks are provided in Appendix B. (c) Regular cubic-lattice packing of $m\times m\times m$ identical balls inside the unit ball with $m=3$, and $R=0.618\,R_{\mathrm{max}}$, with the maximal radius $R_{\mathrm{max}}=L/(\sqrt{3}(m-1)+1)\approx 0.22$.
@@ III.2 Validation
Figure 2: Correlation coefficient $C_{q}(\circ)$ for (a) the unit ball and (b) the unit disk ($R=1$). Comparison between Monte Carlo results (MC, blue dots) and analytical expressions (41, 42) (orange curves). Numerical statistics are obtained from $N=10^{6}$ particles, with $D=1$. Green dashed lines indicate the asymptotic behaviors $C_{q}(\circ)\approx 5\sqrt{7/13}/(qR)$ for $d=3$ and $C_{q}(\circ)\approx 4\sqrt{3/5}/(qR)$ for $d=2$ in the limit $q\to\infty$.
@@
Our Monte Carlo method is validated by comparing the correlation coefficients $C_{q}(\circ)$ obtained from simulations with the exact analytical expressions (41, 42) for uniform starting points inside the unit ball ($d=3$) or the unit disk ($d=2$). As shown in Fig. 2(a), the numerical results (blue dots) are in excellent agreement with Eq. (41) (orange curve) across a broad range of $q$, from $10^{-2}$ to $10^{3}$. In the limit $q\to 0$, one gets $C_{q}\to 1$, indicating that the FRT $\tau$ and the BLT $\ell_{\tau}$ become fully correlated. As $q$ increases, $C_{q}(\circ)$ decreases monotonically. In the large reactivity limit $q\to\infty$, a leading-order expansion of Eq. (41) yields the asymptotic behavior $C_{q}(\circ)\propto 1/q$. The same behavior is observed for the unit disk as shown in Fig. 2(b), where the simulation data perfectly match the exact solution given by Eq. (42).
@@
Figure 3: Joint histograms of the FRT $\tau$ and the BLT $\ell_{\tau}$ for $q=10^{-2},10^{-1},10^{0},10^{1}$ obtained by Monte Carlo simulations with $N=10^{6}$ particles, where the starting points are uniform inside the unit ball. Both $\tau$ and $\ell_{\tau}$ are normalized by their respective mean values. Each blue dot represents a random realization of the pair $(\tau,\ell_{\tau})$, whereas their scatter visualizes correlations. The marginal distributions of $\tau/\langle\tau\rangle$ and $\ell_{\tau}/\langle\ell_{\tau}\rangle=q\ell_{\tau}$ are shown above and right of each joint plot.
@@
Beyond the correlation coefficient itself, the joint histogram of $\tau$ and $\ell_{\tau}$ offers a more direct visual representation of their statistical coupling. Figure 3 presents such joint histograms in the case of uniform starting points inside the unit ball at different reactivities. The random variables $\tau$ and $\ell_{\tau}$ are normalized by their mean values given by Eq. (50a) and Eq. (13), respectively. For large $q$, the pairs $(\tau,\ell_{\tau})$ are widely scattered, reflecting weak correlations. In turn, for smaller reactivity, pairs concentrate along a straight line $\ell_{\tau}/\langle\ell_{\tau}\rangle\approx\tau/\langle\tau\rangle$, indicating the strong linear relation between $\tau$ and $\ell_{\tau}$ anticipated from the limit $C_{q}\to 1$ as $q\to 0$. In addition, we present the marginal distributions of $\tau$ and $\ell_{\tau}$ that exhibit similar exponential decay across all $q$. One sees that most random realizations of $(\tau,\ell_{\tau})$ are concentrated near the lower bottom corner of each panel, i.e., in the region where $\tau/\langle\tau\rangle$ and $\ell_{\tau}/\langle\ell_{\tau}\rangle$ are of the order of $1$ (or smaller).